# Jaynes-Cummings model in a unitary fractional-time description

### v1

@@ IV Entanglement between atom-field
The entanglement phenomenon has been discussed since the beginning of quantum mechanics [47]. It describes a scenario in which the quantum states of two or more particles become so correlated that the state of one particle cannot be described independently of the others, even when they are spatially separated. This is traditionally viewed as a manifestation of non-separability [48]. In the context of Bell nonlocality, not all entangled states violate Bell’s inequality, but any state that does violate it must be entangled. Thus, entanglement is a necessary condition for violating Bell’s inequality [49]. This violation reflects deviations in the statistical correlations of quantum states from classical expectations based on local realism [50, 51]. These concepts underscore the nonclassical nature of quantum mechanics and carry profound implications for our understanding of reality.
@@
In this section, we explore the influence of the FTSE on quantum entanglement. Specifically, in Refs. [28, 29], the FTSE is employed to investigate entanglement within the JC model. Here, we analyze the entanglement dynamics based on the unitary dynamics associated with the FTSE, using the Dyson map as previously discussed. To quantify entanglement in pure bipartite states, the von Neumann entropy [52] can be used as a measure. It is defined as:
@@
$\hat{\rho}(t)=\hat{u}_{\alpha}(t)|\psi^{\alpha}(0)\rangle\langle\psi^{\alpha}(0)|\hat{u}_{\alpha}^{\dagger}(t),$ (27)
@@
we can obtain the reduced density matrix $\hat{\rho}_{\text{a}}(t)$ of the atom by tracing over the degrees of freedom of the field, which gives us
$\hat{\rho}_{\text{a}}(t)=\left[\begin{matrix}\rho_{ee}^{\alpha}(t)&\rho_{eg}^{\alpha}(t)\\ [\rho_{eg}^{\alpha}(t)]^{\ast}&\rho_{gg}^{\alpha}(t)\end{matrix}\right],$ (28)
where we use the same initial state of the previous section to obtain the matrix entries $\rho_{ee}^{\alpha}(t)=\sum_{n=0}^{\infty}|A_{e,n}^{\alpha}|^{2}$, $\rho_{gg}^{\alpha}(t)=\sum_{n=0}^{\infty}|A_{g,n}^{\alpha}|^{2}$, and $\rho_{eg}^{\alpha}(t)=\sum_{n=0}^{\infty}A_{e,n+1}^{\alpha}[A_{g,n}^{\alpha}]^{\ast}$. Taking the trace of the reduced density matrix, we verify that $\rho_{ee}^{\alpha}(t)+\rho_{gg}^{\alpha}(t)=1$ having the trace-preserving property as expected. In Fig. 2, we plot the von Neumann entropy evaluated through the Eq. (26) for different values of fractional parameter $\alpha$. We set the coherent parameter $\beta=2$ and the initial Dyson map parameters to be $\kappa_{\alpha}(0)=0$, $\Lambda_{\alpha}(0)=1$ and $\lambda_{\alpha}(0)=0$. For $\alpha=1.00$, in the well-known scenario where the field and atom are initially prepared in a separable pure state, the quantum dynamics for $t>0$ lead to an increase in marginal entropies, resulting in strong entanglement between the field and the atom. A similar behavior is observed for $\alpha=0.75$. However, for $\alpha=0.25$, the entanglement dynamics undergo a significant change. The entropy initially increases but, after some oscillations, it suddenly drops to nearly zero, indicating that the composite system returns to an almost separable state. This phenomenon suggests the occurrence of the sudden birth and death of entanglement, characterized by a rapid increase in entropy followed by a swift decrease to nearly zero.
@@
Figure 2: Time-evolution of the von Neumann entropy for different values of the fractional-order parameter $\alpha$. We plot the usual case corresponding to $\alpha=1.0$ (black solid line), $\alpha=0.75$ (blue dashed line) and $\alpha=0.5$ (orange dotted line). We start with the atom in the excited state and the field in the coherent state with $\beta=2$, and the initial values of Dyson map parameters being $\kappa_{\alpha}(0)=0$, $\Lambda_{\alpha}(0)=1$ and $\lambda_{\alpha}(0)=0$.
V Conclusions
@@
To summarize, we investigate the dynamics of the JC model within the fractional-time scenario, utilizing the Caputo derivative as outlined in Ref. [37]. This methodology leverages the time-dependent non-Hermitian Hamiltonian theory [38, 40], where a time-dependent Dyson map is constructed to link a dynamic Hilbert space to the time evolution governed by a time-dependent non-Hermitian Hamiltonian, which is unitary. Similarly, this approach is applied in the fractional-time scenario, as demonstrated in Ref. [37] for a traceless two-level system with a general non-Hermitian Hamiltonian. This framework addresses the issue of non-unitarity and aligns with quantum mechanical principles. To apply this unitary description to the JC model, represented in an infinite-dimensional space, we decompose it into invariant two-dimensional subspaces. We then examine how the fractional-order parameter $\alpha$ in the unitary dynamics affects the collapse and revival phenomena, which are intrinsic to the quantum nature of the electromagnetic field and absent in classical systems. Additionally, we explore the atom-field entanglement across different fractional-order parameters, revealing the potential for both the emergence and disappearance of entanglement. Entanglement, a quintessential quantum phenomenon with no classical analog, provides valuable insights into quantum mechanics and is of significant interest in quantum optics and quantum information science. Our results give some connections between fractional-time and non-Hermitian quantum mechanics. By providing this analysis, we hope our results may interest the community working on fractional-time and non-Hermitian systems by inspiring the search for their unitary description, which is generally ignored in those approaches.
@@ Acknowledgements
The author thanks Maike A. F. dos Santos for his valuable comments and clarifications that significantly improved the work. Thanks are also due to Bárbara Amaral and the members of the Quantum Foundation Group (IFUSP) for their insightful discussions. Additionally, the author would like to acknowledge financial support from Instituto Serrapilheira, and the Pró-Reitoria de Pesquisa e Inovação (PRPI) from the Universidade de São Paulo (USP) by financial support through the Programa de Estímulo à Supervisão de Pós-Doutorandos por Jovens Pesquisadores.

### v2

@@ V Atom-field Entanglement
The entanglement phenomenon has been discussed since the beginning of quantum mechanics [52]. It describes a scenario in which the quantum states of two or more particles become so correlated that the state of one particle cannot be described independently of the others, even when spatially separated. This is traditionally viewed as a manifestation of non-separability [53]. In Bell nonlocality, not all entangled states violate Bell’s inequality, but any state that does violate it must be entangled. Thus, entanglement is a necessary condition for violating Bell’s inequality [54]. This violation reflects deviations in the statistical correlations of quantum states from classical expectations based on local realism [55, 56]. These concepts underscore the nonclassical nature of quantum mechanics and carry profound implications for our understanding of reality.
@@
We explore the influence of the FTSE on quantum entanglement. Previous studies [27, 28] have explored entanglement within the JC model using FTSE. Here, we analyze the entanglement dynamics based on the unitary framework associated with the FTSE, using the TD Dyson map as previously discussed. To quantify entanglement in pure bipartite states, we employ the von Neumann entropy [57], a well-established measure of entanglement, defined as:
@@
The composite state given in Eq. (26) corresponds to the density operator
$\hat{\rho}^{\alpha}(t)=|\psi_{\alpha}(t)\rangle\langle\psi_{\alpha}(t)|.$ (30)
@@
Performing a partial trace over the field subsystem, we obtain the reduced density matrix of the atom, $\hat{\rho}_{\text{a}}^{\alpha}(t)$, given by
$\hat{\rho}_{\text{a}}^{\alpha}(t)=\left[\begin{matrix}P_{e}^{\alpha}(t)&r_{\alpha}(t)\\ [r_{\alpha}(t)]^{\ast}&P_{g}^{\alpha}(t)\end{matrix}\right],$ (31)
where the diagonal terms, $P_{e}^{\alpha}(t)$ and $P_{g}^{\alpha}(t)$, represent the population of the excited and grounded states, respectively, as discussed earlier. The off-diagonal coherence term is given by $r_{\alpha}(t)=\sum_{n=0}^{\infty}A_{e,n+1}^{\alpha}[A_{g,n}^{\alpha}]^{\ast}$. Note that the trace-preserving property of the reduced density matrix is verified, as $P_{e}^{\alpha}(t)+P_{g}^{\alpha}(t)=1$.
@@
Figure 2: Time evolution of the von Neumann entropy $S_{\alpha}(t)$ for different values of $\alpha$. The horizontal dashed line indicates the upper bound for the von Neumann entropy $S_{\text{max}}=\ln 2\approx 0.693$. It is considered initially the atom in the excited state and the field in the coherent state with $\beta=2$, and the initial values of Dyson map parameters being $\kappa_{\alpha}(0)=0$, $\Lambda_{\alpha}(0)=1$ and $\lambda_{\alpha}(0)=0$, with the atom-field coupling $\mu_{\alpha}=1$. We analyze for $\alpha=1.00$ (a), $\alpha=0.75$ (b), $\alpha=0.60$ (c), and $\alpha=0.50$ (d), $\alpha=0.45$ (e), $\alpha=0.40$ (f), $\alpha=0.30$ (g), and $\alpha=0.25$ (h).
For the density matrix in Eq. (31), we denote the von Neumann entropy by $S_{\alpha}(t)$. Then, from Eq. (29), it follows that
$S_{\alpha}(t)=-\vartheta_{+}^{\alpha}(t)\ln\vartheta_{+}^{\alpha}(t)-\vartheta_{-}^{\alpha}(t)\ln\vartheta_{-}^{\alpha}(t),$ (32)
where the $\vartheta_{\pm}^{\alpha}(t)$ denote the eigenvalues of the reduced density matrix $\hat{\rho}_{\text{a}}^{\alpha}(t)$, given by
$\vartheta_{\pm}(t)=\frac{1}{2}\pm\frac{1}{2}\sqrt{W_{\alpha}^{2}(t)+4|r_{\alpha}(t)|^{2}},$ (33)
with $W_{\alpha}(t)$ being the population inversion expressed in Eq. (IV). To quantify the entanglement dynamics, we numerically evaluate the von Neumann entropy, as defined in Eq. (32), for different values of $\alpha$. The results are depicted in Fig. 2. In the plot, the horizontal dashed line indicates the upper bound for the von Neumann entropy for a two-level system, which is given by $S_{\text{max}}=\ln 2\approx 0.693$. Similar to the previous analysis, we set the atom-field coupling to be $\mu_{\alpha}=1$, the coherent parameter $\beta=2$, and the initial TD Dyson map parameters to be $\kappa_{\alpha}(0)=0$, $\Lambda_{\alpha}(0)=1$ and $\lambda_{\alpha}(0)=0$. The case of $\alpha=1.00$ (Fig. 2a) corresponds to the well-known JC model. Starting with a separable pure state for the field and atom, the quantum dynamics, for $t>0$, lead to increasing marginal entropy and strong entanglement between the field and the atom [58, 59]. In this case, the entanglement reaches its maximum value at various instants of time, while for $1>\alpha\geq 0$, it approaches the maximum with a slight deviation. For $\alpha=0.75$ (Fig. 2b) the entanglement reaches its maximum value only at the beginning of evolution and remains oscillating below the upper bound. When $\alpha=0.60$ (Fig. 2c), the entanglement oscillates without reaching its maximum value within the plotted time. For $\alpha=0.50$ (Fig. 2d), the entanglement exhibits a more periodic pattern. It is characterized by an initial rise in entropy, subsequent oscillations, and a quick return to a nearly separable state, suggesting entanglement birth and death cycles. As observed in the case of the population inversion, for $\alpha<0.50$, large oscillations start to appear in entanglement entropy. This is illustrated for $\alpha=0.45$ (Fig. 2e), $\alpha=0.40$ (Fig. 2e), $\alpha=0.30$ (Fig. 2g), and $\alpha=0.25$ (Fig. 2h). Despite large fluctuations, the behavior for $\alpha=0.25$ is quasi-periodic.
VI Conclusions
@@
To summarize, we describe the dynamics of the JC model within a fractional-time to a unitary framework, employing the Caputo derivative as outlined in Ref. [32] for a traceless two-level system with a general non-Hermitian Hamiltonian. This framework links the FTSE to the TD non-Hermitian Hamiltonian theory [33, 34], where a TD Dyson map is constructed to define a dynamic Hilbert space with a modified inner product with respect to which the fractional time evolution is unitary. It circumvents the issue of nonunitarity and aligns with quantum mechanical postulates. To apply the results from Ref. [32] to the infinite-dimensional JC model, we decompose it into dynamical invariant two-dimensional subspaces. We then examine how the fractional-order parameters within the unitary dynamics affect the collapse and revival phenomenon, which are intrinsic to the quantum nature of the electromagnetic field and absent in classical systems. We also explore the atom-field entanglement across different fractional-order parameters, revealing the potential for both the birth and death of entanglement. Entanglement, a genuine quantum phenomenon with no classical analog, provides valuable insights into quantum mechanics and is of significant interest in quantum optics and quantum information science. As a potential application, the population inversion dynamics described by Eq. (IV) could be emulated using a reverse engineering approach, as presented in Ref. [60]. This approach involves designing a time-dependent control field within the standard JC model to achieve a specific target population inversion profile. As a further idea, this treatment presented here also may be applied to explore the $\kappa$-deformed JC model, where the JC interaction is described by a time-independent non-Hermitian Hamiltonian due to the deformation parameter [61]. We hope our analysis stimulates further research on unitary formulations for fractional-time and non-Hermitian systems, potentially leading to new insights into fundamental physics.
@@ Acknowledgements
The author thanks Maike A. F. dos Santos, Antonio S. M. de Castro, and Thiago T. Tsutsui for their valuable comments and clarifications that significantly improved the work. Thanks are also due to Bárbara Amaral and the members of the Quantum Foundation Group (IFUSP) for their insightful discussions. Additionally, the author would like to acknowledge financial support from Instituto Serrapilheira, and the Pró-Reitoria de Pesquisa e Inovação (PRPI) from the Universidade de São Paulo (USP) by financial support through the Programa de Estímulo à Supervisão de Pós-Doutorandos por Jovens Pesquisadores.